Two cyclist race against the clock in a 50.0km cross country route. Cyclist a constant speed of 50.0km/h. Cyclist b started 15.0 min after cyclist a but manages to catch up at the finish line. What is the speed of cyclist b assuming that his speed is constant?
1
Expert's answer
2020-10-19T13:20:11-0400
"s_1=s_2\\vt=V \left( t-\frac{15}{60} \right)\\\frac{50}{50}50=V \left(\frac{50}{50}-\frac{15}{60} \right)\\"

"v=66.7\frac{km}{h}"